OpenStudy (anonymous):

Jessica is converting 60 feet per hour into yards per hour. Which conversion factor should she use? A.60 min/1 hr B.1hr/601min C.3ft/1yd D.1yd/3ft

OpenStudy (anonymous):

Will Give medal

OpenStudy (anonymous):

It isn't A or B

OpenStudy (anonymous):

Is it C

OpenStudy (lasttccasey):

Well put the units over each other to get a clear view,\[60 \frac{ ft }{ hr }*\frac{1* yd }{3* ft } \rightarrow 20\frac{ yd }{ hr }\]when you multiply, the feet will cancel leaving you with yd/hr. You have to multiple so that the units you don't want will cancel.

OpenStudy (anonymous):

So it would be D because 1 yd/3ft

OpenStudy (lasttccasey):

Yep, that is the correct answer!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Can you help with 4 more?

OpenStudy (lasttccasey):

I'm sorry I have to leave. Good luck!

OpenStudy (anonymous):

Ok. Thanks anyways!